A researcher examined a random sample of 150 homes in a small city and found that 24 had solar panels installed on their roofs. Use the sample to construct a 99% confidence interval for the proportion of all homes in the city that have solar panels installed on their roofs.

What is the sample proportion? Round to two decimal places.
Answer #1

Solution :

Given that,

n = 150

x = 24

Point estimate = sample proportion = \hat p = x / n = 0.16

1 - \hat p = 0.84

At 99% confidence level the z is ,

\alpha = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.576 * (\sqrt((0.16 * 0.84) / 150)

= 0.077

A 99% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.16 - 0.077 < p < 0.16 + 0.077

0.083 < p < 0.237

A 99% confidence interval for the proportion of all homes in the city that have solar panels installed on their roofs. ( 0.083 , 0.237 )